Using a standard deck of 52 cards and two six-sided dice, what is the probability that you pick or roll the number 2? Exclude the possibility of picking and rolling the number 2.
Answer #1

Number of 2s in a deck of cards = 4

P(picking a 2) = 4/52 = 1/13

P(rolling the number 2 when 2 dice are rolled) = 1 - P(not rolling number 2 on both dice)

= 1 - 5/6 x 5/6

= 11/36

P(picking and rolling number 2) = 1/13 x 11/36

P(you pick or roll the number 2) = 1/13 + 11/36 - (1/13 x 11/36)

= 0.3590